Based on the following information, which of these statements is false?

   Time Period 1st Quarter    2nd Quarter    3rd Quarter

Beginning finished goods (units)    5,000 4,000    12,000

Projected sales units    50,000 40,000 120,000

Desired ending finished goods (units)    4,000 12,000 15,000

Direct labor completes 2 units per hour at $8 per hour

Which of the following statements is false:

Direct labor hours for the 3rd quarter will be 66,000.

Direct labor cost for the 4th quarter cannot be determined.

150,000 sales units are projected for the 4th quarter.

Direct labor cost for the 2nd quarter will be $192,000.
Answer #1

Units to be produced in 3rd quarter = Unit sales + Ending inventory units - Beginning inventory units

= 120,000 + 15,000 - 12,000

= 123,000

Labor hours required = 123,000 / 2 = 61,500

- Direct labor cost for the 4th quarter cannot be determined due to the lack of information.

- 150,000 sales units are projected for the 4th quarter as it is seen that the ending inventory of a quarter is equal to 10% of the projected sales of the next quarter. 4th quarter sales = 15,000 ending inventory of quarter 3 / 10% = 150,000

- Units to be produced in 2nd quarter = Unit sales + Ending inventory units - Beginning inventory units

= 40,000 + 12,000 - 4,000

= 48,000

Labor hours required = 48,000 / 2 = 24,000

Labor cost = 24,000 hours * $8 = $192,000

The answer is Direct labor hours for the 3rd quarter will be 66,000.